OpenStudy (anonymous):

The change in water level of a lake is modeled by a polynomial function, W(x). Describe how to find the x-intercepts of W(x) and how to construct a rough graph of W(x) so that the Parks Department can predict when there will be no change in the water level. You may create a sample polynomial of degree 3 or higher to use in your explanations.

OpenStudy (anonymous):

@Michele_Laino Ok

OpenStudy (michele_laino):

a general formula for a polynomial of third degree can be this: \[\Large W\left( x \right) = A{x^3} + B{x^2} + Cx + D\] where A, B, C, and D are real coefficients

OpenStudy (anonymous):

Ok. Know what do I do?

OpenStudy (michele_laino):

in order to determine those coefficients, we need to know some data about the water level of your lake. Do you have a table which collects the water level of the lake in function of time?

OpenStudy (anonymous):

No. That was the whole question. That's why I was confused :(

OpenStudy (michele_laino):

if we draw a graph in which time x is along the horizontal axis, and the water level W(x) is along the vertical axis, then the x-intercept, is the month at which W(x)=0

OpenStudy (michele_laino):

So a way to establish the x-intercept, is to record at which month, the water level W(x)=0

OpenStudy (anonymous):

Is that the answer?

OpenStudy (michele_laino):

the complete answer should report the values of each coefficient A, B, C, and D, and the followed procedure to get those values. Nevertheless, without any data about the water level of the lake, we are not able to evaluate those coefficients

OpenStudy (anonymous):

what about the last part? Should we create an imaginary problem with a degree higher than 3?

OpenStudy (michele_laino):

Using a polynomial whose degree is greater than 3, involves many real coefficients whixh have to be determined. For example, if we conjecture a polynomial of degree 4, then we can write: \[\Large W\left( x \right) = {A_1}{x^4} + {A_2}{x^3} + {A_3}{x^2} + {A_4}x + {A_5}\] As you can see, now we have to determine, by using our experimental observations, 5 real coefficients, namely: \[\Large {A_1},\;{A_2},\;{A_3},\;{A_4},\;{A_5}\]

OpenStudy (michele_laino):

which*

OpenStudy (anonymous):

ahhh I understand! Ok so thats the final answer, correct?

OpenStudy (michele_laino):

yes! You can write this: "We can establish the values of each coefficients of our sample polynomial function for W(x), by experimental observations about the water level W(x) as function of time or as function of the months of the year"

OpenStudy (anonymous):

Thank-you:)

OpenStudy (michele_laino):

:)

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

I think that better is if we make a drawing of that function: \[\Large T\left( x \right) = {\left( {x - 4} \right)^3} + 6\] That function is represented by a cubic parabola

OpenStudy (anonymous):

I agree

OpenStudy (michele_laino):

here is the corresponding graph:

OpenStudy (michele_laino):

as we can see the turning point is at x=4. At x=4 the temperature is: \[\Large T\left( 4 \right) = {\left( {4 - 4} \right)^3} + 6 = 0 + 6 = 6\]

OpenStudy (michele_laino):

yes! I think so!

OpenStudy (michele_laino):

The requested experimental procedure, which can be used in order to find that turning point, can be this: "We substitute many values for the x variable, when a change in x, produces little change in T(x), then we are close to that turning point"

OpenStudy (anonymous):

I honestly wanna say thank-you for all of your help::)

OpenStudy (michele_laino):

:)

OpenStudy (anonymous):

Sorry, I have one last question!

OpenStudy (michele_laino):

I'm pondering...

OpenStudy (michele_laino):

as stated in the previous exercise, a turning point can be like a vertex of a parabola, or a quadratic function

OpenStudy (michele_laino):

so, Tucker and Karly are both correct, if they refer to a graph like this: